From GlueXWiki
Jump to: navigation, search

April 24, 2014 FDC meeting


  1. Gas system [1],[2] (Dave, Beni)
  2. CDC status (Beni, Mike)
  3. DAQ [3], [4] (Beni)
  4. Electronics (Chris, Nick)
  5. Experimental Safety Assessment Document HallD_ESAD (Lubomir, Beni)
  6. FDC cathode scan [5] (Bill, Lubomir)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Eugene, Bill, Mike, Dave, Beni, Chris, Nick, Simon, and Lubomir (JLab).

Gas system

- Dave: The pressure tests of the gas system started, we found one bad flow controller that Bill is going to send back for repair. Scott connected the CDC and the FDC mixing vessels so that, without one flow controller, we still can supply gas, the same mixture, for both detectors. The fact that CDC and FDC are going to use slightly different gas mixtures is not important in the moment (Curtis).

- It turned out that we have industrial gas (both Ar and CO2) ordered and used now for the tests of the gas system. An option would be to put filters and continue using it till we get the right gas. Eugene: we should stop using it right away, or when we are done with the pressure tests. Dave prepared documents (attached above) for the flushing of the gas system and the detectors, but it should wait till we get the right gas. We have not decided at this meeting if we need gas filters in addition to what is already in the system, but Bill is looking into different options.

CDC status

- Mike: the CDC cooling has been installed but the temperature oscillates (Nick) by about 5-6degC. Bill thinks there are easy ways to fix this and he is looking into this. Before the cooling system was installed, Nick measured the temperature on the regulators and it was ~60degC which (according to Fernando) means about 100degC on the silicon. This is not acceptable and the LV should not be kept for more than 1 hour without cooling.


- Cody is modifying the fADC125 firmware to make the beginning and the width of the time window parameters. The numbering of the fADC channels now starts from 0. Beni wants to make sure all this works before he starts testing the CDC.

- Now every detector has its own account to be used with the DAQ. So far Beni was running the DAQ on his own account.

- Beni was reading the FDC strips for more than a week. The results are linked above: two plots for the up- and down-stream cathodes in each a column represents one package. Since there's no trigger even 1Hz noise (similar to the cosmics rate) is visible. Therefore we should try using a trigger, the easiest for now will be to use the scintillator paddles below the magnet (Eugene).


- Some of the cables with bad strip channels that were tested with a scope and looked OK, have changed in the latest Beni's results. This may indicate something is wrong with the 3M connectors. Since we are sure about this for at least 5-6 connectors, Chris is going to look at them tomorrow.


- Beni and Lubomir worked on the CDC/FDC sections of the Hall D ESAD document, linked above. We discussed the safe pressure limits for the two detectors: should be between 0 and 200 Pa and under-pressure is even more dangerous. Actually CDC has never been operated at more than ~100Pa but if we want to have 3 gas exchanges per day the pressure will be higher than that. Beni: important is to make sure the input bubbler can limit the pressure in the detectors. For the FDC the pressure difference between the cells should be <30Pa. LV hardware interlock for the CDC: can be connected to the airflow system directly (like the fan), while the PLS system will measure the temperature and stops the LV.

- Gas sytem text is not ready. Possible hazards: the alcohol in the gas makes it combustible (Bill), high pressure section (above 15 psi).

FDC cathode scan

- Lubomir showed new results (picture linked above): the gaps are now NOT in between the foils, or strip #48 and #144, BUT at #44 and #149. The latter are the places where the Kapton tape starts/ends. The pitch in the side foils is 5.01mm while 5.00mm in the middle foil. As a result of these changes the deviations from the nominal are much smaller than before, at most +/-100 microns. Now both, the right and the left edges are used in the analyses. Somehow the right edge shows smaller variations. It is not clear at all why in the middle foil the pitch didn't change after the tensioning.

- When looking at all measured points along the gaps between the strips one can recognize that some areas are much more scattered than the rest, which indicates problems with the measurements: either bad contrast due to the coper or Kapton tape, or some other systematics. Bill will look if we can re-scan the cathode at least partially with the JLab CMM machines.